Novak Inc. has issued three types of debt on January 1, 2020, the start of the company’s fiscal year.

(a) $11 million, 10-year, 13% unsecured bonds, interest payable quarterly. Bonds were priced to yield 11%.
(b) $27 million par of 10-year, zero-coupon bonds at a price to yield 11% per year.
(c) $20 million, 10-year, 9% mortgage bonds, interest payable annually to yield 11%.


Prepare a schedule that identifies the following items for each bond: (1) maturity value, (2) number of interest periods over life of bond, (3) stated rate per each interest period, (4) effective-interest rate per each interest period, (5) payment amount per period, and (6) present value of bonds at date of issue. (Round stated and effective rate per period to 2 decimal places, e.g. 10.25%. Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971.)

Unsecured
Bonds

Zero-Coupon
Bonds

Mortgage
Bonds

(1) Maturity value $ $ $
(2) Number of interest periods
(3) Stated rate per period % % %
(4) Effective rate per period % % %
(5) Payment amount per period $ $ $
(6) Present value $ $ $
Answer #1

Answer:

Unsecured Bonds Zero-Coupon Bonds Mortgage Bonds
1 Maturity value $11,000,000 $27,000,000 $20,000,000
3 Stated rate per period 3.25% 0 9%
4 Effective rate per period 2.75% 0 11%
5 Payment amount per period $357,500 0 $1,977,800.00
6
Present Value
$12,324,295.56 $9,508,980.93 $17,644,307.20

For Unsecured Bonds:

Maturity Value = $11,000,000

Number of periods = 10 * 4 = 40 (since interest is payable quarterly)

Stated Rate per period = 13%/4 = 3.25%

Effective Rate per period = 11%/4 = 2.75%

Payment Amount per period = 3.25% * 11,000,000 = $357,500

Present Value =PV(2.75%,40,357500,11000000) = $12,324,295.56

For Zero Coupon Bond

Maturity Value = $27,000,000

Since it is a zero coupon bond, there will be no interest paid.

Thus, no of interest periods = stated rate per period = effective rate per period = payment amount per period = 0

Present Value = 27,000,000 /(1 + 11%)10 = $9,508,980.93

For Mortgage Bonds

Maturity Value = $20,000,000

Number of periods = 10

Stated Rate per period = 9%

Effective Rate per period = 11%

Payment Amount per period = 9% * 20,000,000= $1,800,000

Present Value =PV(11%,10,1800000,20000000) = $17,644,307.20